OpenStudy (anonymous):

Hi ! I am an 8th grade student, I skipped a year in math, and I just wanted to make sure my equation simplifieng for slope is correct. -3x + 2y = 8. Now I know we are using the slope-intercept method, and to do that we have to isolate the "y" so I plussed 3 on both sides of the equation, then we have, 2y = 3x + 8, but "y" isn't by itself yet so we have to divide both sides of the equation by 2 and I got , y = 3x + 4, so my slope is 3 and my y-intercept is 4 right ?

OpenStudy (anonymous):

"but "y" isn't by itself yet so we have to divide both sides of the equation by 2" you need to divide everything on the right by 2, not just the 8

OpenStudy (anonymous):

I did though, I devided by 2.

OpenStudy (anonymous):

\[\frac{1}{2}(3x+8)=\frac{3}{2}x+4\]

OpenStudy (anonymous):

that is what i mean by "everything" you must divide \(3x\) by 2 as well as divide 8 bu 2

OpenStudy (anonymous):

? No , I just need slope lol, not that . I ended up with a final slope equation of y = 3x + 4

OpenStudy (anonymous):

I wasn't taught that way yet though, just to find the second point on the graph using rise/run.

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

you are right to this line \[2y = 3x + 8\]

OpenStudy (anonymous):

okay ! ready

OpenStudy (anonymous):

you want to solve this for \(y\) and you said "divide by 2" right?

OpenStudy (anonymous):

when you divide both sides of an equation by 2 you must divide each term by 2

OpenStudy (anonymous):

No, no ,no lol sorry if I was un clear, I just want to isolate y (get it by itself )

OpenStudy (anonymous):

yes i know that is what you are trying to do

OpenStudy (anonymous):

and you do that by dividing both sides by 2

OpenStudy (anonymous):

when you divide \(2y\) by 2 you get \(y\)

OpenStudy (anonymous):

so we put everything on the right side of the equation over 2

OpenStudy (anonymous):

when you divide \(3x+8\) by 2 you get \(\frac{3}{2}x+4\)

OpenStudy (anonymous):

you cannot just divide the constant 8 by 2 and leave the \(3x\) alone, you have to divide \(3x\) by 2 also

OpenStudy (anonymous):

no, just 3/1

OpenStudy (anonymous):

for slope

OpenStudy (anonymous):

your slope is therefore \(\frac{3}{2}\) not 3

OpenStudy (anonymous):

so.... 3/2 x wold be my slope ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

gotta run, good luck http://en.wikipedia.org/wiki/Beit_Jala

OpenStudy (anonymous):

okay thanx !